2x – 5y = -1
x – 3y = -3

Answer by sophxmai(62)

2x–5y=-1
x–3y=-3

You first want to get rid of one of the variables.
There are two different ways of doing this.
The first way is to solve the 2nd equation for x, and then plug it into the 1st equation to solve for y.
The second way is to multiply the 2nd equation by 2, and then subtract this new equation from the 1st equation so that x cancels out. Left with y, you solve for it.

The first way.
2x–5y=-1
x–3y=-3

Solve the 2nd equation for x
x-3y=-3
x=3y-3

Substitute x into the first equation
2x-5y=-1
2(3y-3)-5y=-1

Simplify and solve for y
2(3y-3)-5y=-1
6y-6-5y=-1
y=5

Plug y back into an equation with x in it.
x=3y-3
x=3(5)-3
x=15-3
x=12

The second way.
2x–5y=-1
x–3y=-3

Multiply the 2nd equation by 2
x–3y=-3
(x-3y=-3)x2
2x-6y=-6

Subtract the new equation from the first equation.
2x–5y=-1
2x-6y=-6
--------------
y=5

Plug y into an equation with x in it.
x–3y=-3
x-3(5)=-3
x-15=-3
x=12